OpenStudy (anonymous):

5 sqrt 11 ( sqrt 3 -3 sqrt 2) i know the answer =41.6 but no idea how to get it. can someone show me how

OpenStudy (anonymous):

\[5\sqrt{11}(\sqrt{3}-3\sqrt{2})\]

OpenStudy (phi):

with a calculator. first do sqr(3)

OpenStudy (phi):

then - 3 times sqr(2) what do you get ?

OpenStudy (anonymous):

4.24

OpenStudy (phi):

can you do it step by step? first what do you get for sqr(3) ?

OpenStudy (anonymous):

1.73

OpenStudy (phi):

now while that number is still in the calculator type - (minus sign) 3 * (times sign) sqr 2 = or enter

OpenStudy (anonymous):

i got -2.51

OpenStudy (phi):

yes. now multiply that by 5 and then multiply that number by sqr(11) you get -41.6 (there are more digits by the way)

OpenStudy (anonymous):

i know lol im just to lazy to write them all out. but thank youc:

OpenStudy (phi):

Ok, just so long as you know. (sometimes people set their calculator to show only a few digits and don't realize the answer is not accurate.)

OpenStudy (phi):

We are getting -41.6 which is different from what you posted up above. I hope that is a typo.

OpenStudy (phi):

you should practice doing the calculation all on one line in your calculator.

OpenStudy (anonymous):

yeah its a typo and okay